OpenStudy (anonymous):

I wil medal and fan for help!~ Ratio ~ Carlos biked 24 miles and ran 6 miles. Write the ratio of the distance Carlos biked to the total distance. 6 to 29 3 to 12 30 to 24 4 to 5

OpenStudy (mertsj):

What is the total distance?

OpenStudy (anonymous):

30 miles

OpenStudy (mertsj):

how far did he bike?

OpenStudy (anonymous):

24

OpenStudy (mertsj):

Then write the fraction with 24 for the numerator and 30 for the denominator. That is the ratio. Reduce it.

OpenStudy (anonymous):

Biked: 24 Ran: 6 Ratio is like a fraction. The problem says ratio of BIKED TO TOTAL DISTANCE. 24+6= 30 24/30 = 4/5 So its D